OpenStudy (anonymous):

2 7/9 divided by 1/4

OpenStudy (anonymous):

\[2\frac{ 7 }{ 9 } \div \frac{ 1 }{ 4 }\]

OpenStudy (aaalive):

give me a sec

OpenStudy (aaalive):

nvm

OpenStudy (anonymous):

Where do you think you will start?

OpenStudy (triciaal):

@Savannah2816 do you know how to change a mixed number into a fraction? to divide by a fraction you invert the fraction and multiply

TheSmartOne (thesmartone):

\(\sf \Large a\frac{b}{c} = \frac{ac+b}{c}\)

OpenStudy (triciaal):